OpenStudy (anonymous):

What is the slope of(1,7) and (3,10)

OpenStudy (amistre64):

subtract one point from the other, then stack y/x and reduce to your hearts desire

OpenStudy (anonymous):

slope = rise / run m = y / x m = (y_2 - y_1) / (x_2 - x_1) m = (10 - 7) / (3 - 1_ m = 3/2